find the center and radius of the circle given by:
x^2+10x+y^2+12y=-12

Answer by scott8148(6628)

complete the squares for x and y

adding (10/2)^2 and (12/2)^2 __ x^2+10x+25+y^2+12y+36=-12+25+36

(x+5)^2+(y+6)^2=7^2

the circle is centered at (-5,-6) with a radius of 7